# A non-Hermitian loop for a quantum measurement

### v1

@@ Abstract
The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$) and with a standard deviation $\sigma=1$. Fig.1 (b) shows the probabilities of having spin up or down (dashed and full lines respectively) as a function of time. The meter acts around $t=5$ amplifying the spin down and reducing the spin up components leading to a situation where the superposition vanishes, the collapse. After the interaction with the meter vanishes the usual time-evolution leads again to a superposition state.
While this scheme may be adjusted to obtain a desired outcome it leaves, however, no much insight on the process of the measurement itself nor the ultimate mechanism leading to the collapse. One may also argue that substantial fine tuning of the non-Hermitian Hamiltonian is necessary to mimic a quantum collapse  [^3] Note that in this case the Hamiltonian describes a loop in parameter space enclosing a vanishing area..
@@
For simplicity, in what follows we will focus on energy measurements A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions. Furthermore, the mathematical substrate for such phenomenon is likely to break many of the usual rules to serve as a proper backdrop for the cut. Within the realm of non-Hermitian systems, one of the most prominent sources of unusual and counter-intuitive physics are the so called exceptional points Kato (1995); Heiss (2012). These are points in parameter space where the eigenvalues and eigenvectors coalesce (see the scheme in Fig.1 (c) illustrating the Riemann surfaces close to an exceptional point), and have been the focus of intense research  Dembowski et al. (2001, 2004); Stehmann et al. (2004); Hu et al. (2017); Miri and Alù (2019). The coalescence at an exceptional point needs to be distinguished from a usual Hermitian degeneracy where a full set of eigenstates is kept. The key difference being the appearance of defectiveness Foa Torres (2019), the lack of a full set of eigenstates. The dependence of the eigenvalues on the parameters around exceptional points is typically non-analytic Heiss (2012).
@@
Although for a $2\times 2$ system non-Hermitian system, the adiabatic solutions are expected to exchange when encircling an exceptional point Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012), this conversion has been predicted to occur in an asymmetric way Uzdin et al. (2011); Berry and Uzdin (2011), the so called chiral state or mode conversion Hassan et al. (2017a) because of non-adiabatic transitions. This phenomenon manifests in a two-state system during the encirclement of an exceptional point within the parameter space. In the realm of Hermitian systems, the adiabatic theorem posits that if a system is initially prepared in one of its eigenstates and the Hamiltonian parameters are gradually altered along a closed trajectory, the system will revert to its initial state, albeit with an additional geometric phase Berry (1984). In contrast, in non-Hermitian systems, a clockwise encirclement around an exceptional point invariably results in the system settling into one of the eigenstates (deemed the preferred state). On the other hand, an anticlockwise journey consistently leads to the emergence of the second eigenstate, as depicted in Fig.1 (d). This intriguing behavior has been observed across various platforms, including waveguides Doppler et al. (2016), and optomechanical systems Xu et al. (2016). At a fundamental level, this state conversion is associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al. (2015).
@@
Remarkably, the dynamics of the instantaneous eigenstates’ weights in non-Hermitian systems diverge significantly from those in Hermitian counterparts. The weights can exhibit extreme fluctuations Berry and Uzdin (2011), and transitions can occur even under slow dynamics, challenging the conventional understanding of adiabaticity Doppler et al. (2016); Berry and Uzdin (2011). Ref. Hassan et al. (2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting and has shown that the state conversion is robust to deformations of the contour if the initial and final points are kept Hassan et al. (2017a). Recent findings further reveal that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al. (2017b); Nasari et al. (2022), thus broadening the phenomenon’s observed scope.
@@
In our present context, chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system. For this, the interaction with the meter should lead to exceptional points in parameter space. The loop starts (and ends) with the Hamiltonian being ${\cal H}_{0}$ which is Hermitian. Right at the beginning of the loop, a crucial decision must be made regarding the direction in which the loop will proceed. Switching between clockwise and anticlockwise loops changes the outcome or the preferred state. Upon the loop’s completion, when the interaction ceases and the Hamiltonian reverts to its original form ${\cal H}_{0}$, the initial superpositions that may have been in the initial state disappear, ushering in a state of certainty. This transition marks a pivotal moment in the dynamics of the system, where the ephemeral nature of quantum superpositions gives way to definitive outcomes, encapsulating the essence of quantum measurement and evolution.
The probability with which the meter chooses between the clockwise and anticlockwise loops should depend on the initial state and follow Born’s rule if the usual quantum mechanics predictions are to be preserved. This renders the dynamical equation not only non-Hermitian but also non-linear. Another point is that, for a sufficiently slow variation of the parameters, one would not need to fine tune the loop to encircle the exceptional point as the conversion has been shown to occur anyway Hassan et al. (2017b); Nasari et al. (2022).
Collapse when there is no measurement apparatus.– Collapse is known to occur even without a measurement apparatus. When there is a measurement apparatus it seems reasonable to model it as an additional non-Hermitian term that acts during the time of the measurement. But how to account for exceptional points in other cases, as when many body interactions are present? A possible path for an answer could be the proliferation of exceptional points reported in Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022).
We underline the fact that any interaction that vanishes after some time (this is, that does not modify permanently the Hamiltonian of the system of interest) adheres to assumption 2. This compliance allows for the autonomy of loops within the parameter space, enabling for spontaneous collapse, or collapse without a meter, to occur. In this case, one should just rather view collapse as emerging from the interaction with other system, encoded in an effective way in the loop in parameter space.
Classical simulators for state vector collapse.– As a corollary, and perhaps most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits and many others. Indeed many existing experiments could be considered as first steps towards classical realizations/emulations of quantum measurements Dembowski et al. (2001, 2004); Doppler et al. (2016). The role of time could be played by a space coordinate along which the non-Hermitian term can be modulated as done in photonic systems Rechtsman et al. (2013).
@@
Final remarks.– We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation. A quantum measurement is seen as resulting from the interaction with other degrees of freedom introducing a non-Hermitian term to an otherwise Hermitian Hamiltonian. The nature of this interaction, having a finite lifespan, motivates using a time-dependent Hamiltonian which, throughout the interaction, describes a closed loop in parameter space. As a possible mechanism we signal the phenomenon called chiral state conversion, which is enabled by exceptional points and leads to the required destruction of superposition states. For a $2\times 2$ system, the outcome can be switched by changing the encircling of the loop from clockwise to anticlockwise. The probability of taking one option or the other, which amounts to choosing how the action of the meter takes place, needs to be given by the initial state so as to recover the usual predictions of quantum mechanics. This makes the dynamic equation also non-linear, which is consistent with other studies Mertens et al. (2021).
@@
Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom. Another mechanism that may come handy is the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018), which essentially leads to an anomalous localization at one edge of a pristine lattice. In some cases one might tune the non-Hermitian term to select a few or even a single eigenstate. There has been a surge of interest in this phenomenon, which has been shown to have a topological origin Okuma et al. (2020); Zhang et al. (2020), see for example Okuma and Sato (2023); Bergholtz et al. (2021); Lin et al. (2023) and references therein.
@@
We hope that our discussion may trigger a fresh look to this enduring enigma. Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
Acknowledgements.
@@
L.E.F.F.T. thanks Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C. Cormick for useful comments. L.E.F.F.T. acknowledges financial support by ANID FONDECYT (Chile) through grant 1211038, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No. 873028 (HYDROTRONICS Project). ICN2 is funded by the CERCA programme / Generalitat de Catalunya, and is supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN / AEI / 10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya.

### v2

@@ III Possible mechanisms
The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$, represented with a vertical dashed line in Fig.1 (b)) and with a standard deviation $\sigma=1$. Specifically, we take $\lambda(t)=\exp(-(t-t_{0})^{2}/(2\sigma^{2}))/(\sigma\sqrt{2\pi})$. The state vector $|\Psi(t)\rangle$ evolves under $\mathcal{H}$ according to the linear Schrödinger equation as outlined in Section II. The probabilities displayed in Fig.1 (b) for having spin up (blue continuous line) and spin down (orange dashed line) as a function of time (the initial state is a spin down state) are then calculated from the components of $|\Psi(t)\rangle$ – whose projections onto the spin basis states are treated as relative probability amplitudes – by squaring their magnitudes and normalizing by the total instantaneous norm $\langle\Psi(t)|\Psi(t)\rangle$. The meter acts around $t=5$ amplifying the spin up and reducing the spin down components leading to a situation where the superposition vanishes, the collapse. After the interaction with the meter vanishes the usual time-evolution leads again to a superposition state.
Although these models show how non-Hermitian Hamiltonians can amplify a chosen state, the need to construct the interaction specifically to favor a predetermined outcome perhaps offers less insight into the measurement process itself. One might desire a mechanism where state selection emerges more dynamically from the system’s response to control or interaction. This motivates our exploration of an alternative framework centered on dynamically induced collapse.
@@
For simplicity, in what follows we will focus on energy measurements. A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions. Furthermore, the mathematical substrate for such phenomenon is likely to break many of the usual rules to serve as a proper backdrop for the cut. Within the realm of non-Hermitian systems, one of the most prominent sources of unusual and counter-intuitive physics are the so called exceptional points Kato (1995); Heiss (2012). These are points in parameter space where the eigenvalues and eigenvectors coalesce (see the scheme in Fig.1 (c) illustrating the Riemann surfaces close to an exceptional point), and have been the focus of intense research  Dembowski et al. (2001, 2004); Miri and Alù (2019). The coalescence at an exceptional point needs to be distinguished from a usual Hermitian degeneracy where a full set of eigenstates is kept. The key difference being the appearance of defectiveness Foa Torres (2019), the lack of a full set of eigenstates. The dependence of the eigenvalues on the parameters around exceptional points is typically non-analytic Heiss (2012).
@@
Although for a $2\times 2$ system non-Hermitian system, the adiabatic solutions are expected to exchange when encircling an exceptional point Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012), this conversion has been predicted to occur in an asymmetric way Uzdin et al. (2011); Berry and Uzdin (2011), the so called chiral state or mode conversion Hassan et al. (2017a) because of non-adiabatic transitions. This phenomenon manifests in a two-state system during the encirclement of an exceptional point within the parameter space. In the realm of Hermitian systems, the adiabatic theorem posits that if a system is initially prepared in one of its eigenstates and the Hamiltonian parameters are gradually altered along a closed trajectory, the system will revert to its initial state, albeit with an additional geometric phase Berry (1984). In contrast, in non-Hermitian systems, a clockwise encirclement around an exceptional point invariably results in the system settling into one of the eigenstates (deemed the preferred state). On the other hand, an anticlockwise journey consistently leads to the emergence of the second eigenstate, as depicted in Fig.1 (d). This intriguing behavior has been observed across various platforms, including waveguides Doppler et al. (2016), and optomechanical systems Xu et al. (2016). At a fundamental level, this state conversion has been associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al. (2015).
@@
Remarkably, the dynamics of the instantaneous eigenstates’ weights in non-Hermitian systems diverge significantly from those in Hermitian counterparts. The weights can exhibit extreme fluctuations Berry and Uzdin (2011), and transitions can occur even under slow dynamics, challenging the conventional understanding of adiabaticity Doppler et al. (2016); Berry and Uzdin (2011). Ref. Hassan et al. (2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting. Recent papers indicate that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al. (2017b); Nasari et al. (2022), thus broadening the phenomenon’s observed scope. But the understanding of the phenomenon of state conversion is still far from complete, as recent studies indicate Nye (2023); Nye and Kantartzis (2024); Kumar et al. (2025). Clearly, more research is needed to get a better picture of the time scales needed for it to happen and its robustness.
@@
Chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system. For this, the interaction with the meter should lead to exceptional points in parameter space. The loop starts (and ends) with the Hamiltonian being ${\cal H}_{0}$ which is Hermitian. Right at the beginning of the loop, a crucial decision must be made regarding the direction in which the loop will proceed. Switching between clockwise and anticlockwise loops changes the outcome or the preferred state. Upon the loop’s completion, when the interaction ceases and the Hamiltonian reverts to its original form ${\cal H}_{0}$, the initial superpositions that may have been in the initial state disappear, ushering in a state of certainty. This transition marks a pivotal moment in the dynamics of the system, where the ephemeral nature of quantum superpositions gives way to definitive outcomes, encapsulating the essence of quantum measurement and evolution.
To make the concept of forming a closed loop in parameter space more concrete, let us consider a minimal example which captures the essential physics of EP encirclement. A general form for such a Hamiltonian can be written relative to its parameters $\delta(t)$ and $\gamma(t)$ as:
${\cal H}(t)=\begin{pmatrix}\delta(t)+i\gamma(t)&J\\ J&-\delta(t)-i\gamma(t)\end{pmatrix}$ (5)
Here, $\mathbb{I}$ is the identity matrix and $J$ represents a real coupling strength, and $\delta(t)$ is a time-varying detuning. The term $\gamma(t)$ represents half the time-varying gain/loss difference between the two states. Its significance is central to the non-Hermitian nature of the Hamiltonian during the measurement interaction; it explicitly introduces gain or loss, modeling the strength and character of the system-apparatus coupling that drives the non-unitary evolution. The specific choice and time-dependence of $\gamma(t)$ (along with $\delta(t)$) define the parametric loop in the non-Hermitian parameter space, which is crucial for the proposed chiral state conversion mechanism. The model requires $\gamma(t)$ to be non-zero only during the transient measurement interaction, returning to zero (as does $\delta(t)$) at the beginning ($t=0$) and end ($t=T$) of the loop, ensuring the Hamiltonian reverts to the Hermitian $\mathcal{H}_{0}$. Let the initial and final Hamiltonian be the Hermitian ${\cal H}_{0}$, corresponding to parameters $\delta=0$, and $\gamma=0$:
${\cal H}_{0}=\begin{pmatrix}0&J\\ J&0\end{pmatrix}.$ (6)
The EP occurs at $(\delta_{EP},\gamma_{EP})=(0,J)$ (assuming $J>0$). A closed loop in the $(\delta,\gamma)$ parameter space that starts from ${\cal H}_{0}$ (i.e., starts at $(\delta=0,\gamma=0)$), encircles this EP, and returns to ${\cal H}_{0}$ at time $t=T$ can be explicitly implemented. For example, consider a circular path centered precisely at the EP $(0,J)$ with radius $R=J$. This path passes through the origin and is parameterized during the interval $0\leq t\leq T$ by:
$\delta(t)=J\sin(\Omega t)$ (7)
and
$\gamma(t)=J(1-\cos(\Omega t)),$ (8)
where $\Omega=2\pi/T$ determines the loop duration $T$. Note that at $t=0$ and $t=T$, we have $\delta=0$ and $\gamma=0$, ensuring the Hamiltonian starts and ends at ${\cal H}_{0}$. The time evolution governed by this explicit ${\cal H}(t)$ generates the phenomenon of chiral state conversion Hassan et al. (2017a).
[…]
To connect our model to the probabilistic outcomes of quantum measurements, we must address how the choice of loop direction (chirality) relates to Born’s rule. Our framework proposes that the non-Hermitian interaction drives the system towards a specific eigenstate depending on the chirality (loop direction). However, the model in its current form does not inherently derive the probabilities associated with these outcomes. To recover the predictions of conventional quantum mechanics, we must impose Born’s rule externally: the probability of the interaction selecting a particular chirality (clockwise or anti-clockwise) must be explicitly set to match the squared amplitude of the corresponding component in the system’s state vector just before the measurement interaction begins. The precise physical mechanism responsible for enforcing this probabilistic choice based on the initial state remains an open question not solved by this dynamical model. Requiring the chirality selection probability to depend on the initial state amplitudes inherently introduces a non-linear element into the overall measurement process description, distinguishing it from the purely linear Schrödinger evolution between measurements. This aligns with arguments suggesting that a purely linear dynamics cannot fully replicate Born’s rule  Mertens et al. (2021). Importantly, for a sufficiently slow variation of the parameters, encircling the exceptional point might not be necessary, as the conversion has been signaled to occur anyway Hassan et al. (2017b); Nasari et al. (2022). This robustness may simplify the experimental implementation of our proposed mechanism.
Looking ahead, a crucial question is how the probabilistic nature of quantum measurement emerges from the seemingly deterministic chiral dynamics. We speculate that, as long as the environmental interactions (at the origin of the non-Hermitian term) are random and unbiased, they will lead to a statistical distribution of chirality choices that mirrors the probabilities encoded in the quantum state, hence statistically aligning with Born’s rule and the predictions of conventional quantum mechanics. Exploring this connection between environmental interactions, non-Hermitian dynamics, and the emergence of Born’s rule is a promising direction for future research.
Collapse when there is no measurement apparatus.– The collapse of the wave function can occur even in the absence of a dedicated measurement apparatus. While our model focuses on the case where a non-Hermitian term represents the measurement interaction, extending it to other scenarios, such as those involving many-body interactions, remains an open question. A possible path forward could involve exploring the proliferation of exceptional points Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022). Importantly, any interaction that vanishes after a finite time, leaving the system’s Hamiltonian unchanged, automatically satisfies our assumption 2, allowing for the possibility of spontaneous collapse driven by transient interactions represented as loops in parameter space.
IV Final remarks
@@
We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation. A quantum measurement is seen as resulting from the interaction with other degrees of freedom introducing a non-Hermitian term to an otherwise Hermitian Hamiltonian. The nature of this interaction, having a finite lifespan, motivates using a time-dependent Hamiltonian which, throughout the interaction, describes a closed loop in parameter space. As a possible mechanism, we signal the phenomenon called chiral state conversion, which leads to the destruction of superposition states in favor of one of the eigenstates of the initial Hamiltonian (which, for our purposes, is Hermitian). For a $2\times 2$ system, the outcome can be switched by changing the encircling of the loop from clockwise to anticlockwise. Crucially, to comply with experimental observations, the probability of the interaction selecting a clockwise versus anti-clockwise loop must be imposed according to Born’s rule, determined by the amplitudes of the system’s state vector immediately prior to the interaction. As discussed, incorporating this state-dependent probability selection introduces the necessary non-linearity into the description of the measurement process Mertens et al. (2021).
Conceptually, our analysis utilizing parameter loops around exceptional points offers insight into two key aspects of quantum measurement. Firstly, it provides a dynamical mechanism for state selection based on path topology (chirality). Secondly, and perhaps more fundamentally, the dynamics inherent to the non-Hermitian evolution during the loop naturally establish the eigenstates of the initial Hamiltonian $\mathcal{H}_{0}$ as the unique stable outcomes, offering a potential mechanistic explanation for their privileged status in this context. By signaling how the system is dynamically driven towards these states before one is selected, our framework provides a more complete picture of the collapse process.
@@
Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom, where other mechanisms such as the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018) (see also  Bergholtz et al. (2021); Lin et al. (2023) and references therein) may come handy.
As a corollary and perhaps the most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits, and many others. For the case of photonic waveguide arrays Szameit et al. (2011); Vicencio et al. (2025), the direction of propagation $z$ naturally takes the role of time $t$ Rechtsman et al. (2013), owing to the direct analogy between the paraxial wave equation for the light envelope and the time-dependent Schrödinger equation. The effective Hamiltonian maps onto the structure’s properties: coupling $J(z)$ via waveguide spacing, detuning $\delta(z)$ via refractive index $n(x,y,z)$, and non-Hermitian terms $\gamma(z)$ via engineered gain/loss (e.g., complex $n$ or coupling to lossy elements). Since these parameters can be precisely modulated along $z$ during fabrication (e.g., via laser writing), static waveguide circuits can be created to implement specific time-dependent Hamiltonians, including the parameter loops around exceptional points central to our mechanism. This allows for direct observation of the resulting dynamics, including testing the chiral state selection by comparing structures fabricated with opposite loop winding directions along $z$.
@@
We hope that this may trigger a fresh look at this enduring enigma. Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
V Acknowledgments
@@
We thank Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C. Cormick, Michael Berry, Franck Laloë and Quancheng Liu for useful comments. L.E.F.F.T. acknowledges financial support by ANID FONDECYT (Chile) through grants 1211038 and 1250751, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No. 873028 (HYDROTRONICS Project). ICN2 is funded by the CERCA programme / Generalitat de Catalunya, and is supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN / AEI / 10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya.
VI Author Contributions
Building on their established research partnership, L. E. F. F. T. and S. R. collaborated on this new line of inquiry starting with a question posed by S. R.. For this paper, L. E. F. F. T. developed the novel concept, led the project, and drafted the manuscript. S. R. contributed valuable insights and critical feedback throughout the process. Both authors were involved in refining the ideas and revising the final manuscript.